OpenStudy (anonymous):

Stuck on a factoring question!! it's 8y^3=2y (^ is exponent.) From there I got, 8y^3-2y=0 2y(4y^2-1)=0 I don't know where to go from there!

OpenStudy (anonymous):

first do exponents

OpenStudy (ranga):

a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

then i think it should be easy to do it

OpenStudy (anonymous):

We use difference of squares to get 2y(2y-1)(2y+1).

OpenStudy (dumbcow):

once you have factors equal to 0 split up the factors and solve each separately --> 2y = 0 --> y=0 --> 4y^2 -1 = 0 4y^2 = 1 y^2 = 1/4 y = +-sqrt(1/4) = +-1/2

OpenStudy (ranga):

Are you asked to solve or just factor?

OpenStudy (anonymous):

both

OpenStudy (ranga):

Okay. 2y(2y-1)(2y+1) = 0 y = 0, y = 1/2, y = -1/2

OpenStudy (isaiah.feynman):

Alternatively, you could also do that...but be sure your factoring is not shaky